OpenStudy (keeli_grace):

Which conjunction or disjunction is equivalent to |t – 7| > 12? A. t – 7 > 12 or t – 7 < –12 B. t – 7 < 12 and t – 7 > –12 C. t – 7 < 12 or t – 7 > –12 D. t – 7 > 12 and t – 7 < –12

OpenStudy (keeli_grace):

@Alexandra675 @jbeanz97 @e.mccormick

OpenStudy (e.mccormick):

For |x| = 1, it means x could be one of two things, -1 or 1. So you would solve it by saying: x = 1 or x = -1 The solve each one. When we take the next harder case: |x+a| = b We do: x+a=b or x+a=-b The solve these to: x=b-a or x=-b-a You need to do something along those lines.

OpenStudy (e.mccormick):

With an inequality, there is a little more to it. The answer might be above or below, or it might be between.